# Surface hopping molecular dynamics simulation of ultrafast methyl iodide photodissociation mapped by Coulomb explosion imaging

### v1

@@
[…]
We present a highly efficient method to directly simulate the photodissociation followed by Coulomb explosion of methyl iodide. In order to achieve statistical reliability, more than 40,000 trajectories are calculated on accurate potential energy surfaces of both the neutral molecule and the doubly charged cation. Non-adiabatic effects during photodissociation are treated using a Landau-Zener surface hopping algorithm. The simulation is performed analogous to a recent pump-probe experiment using coincident ion momentum imaging [Ziaee et al., Phys. Chem. Chem. Phys. 2023, 25, $9$]. At large pump-probe delays, the simulated delay-dependent kinetic energy release signals show overall good agreement with the experiment, with two major dissociation channels leading to $\text{I}(^{2}\text{P}_{3/2})$ and $\text{I}^{*}(^{2}\text{P}_{1/2})$ products. At short pump-probe delays, the simulated kinetic energy release shows a clear bifurcation near 12 fs, owing to non-adiabatic transitions through a conical intersection. The developed method is particularly suitable and efficient in simulating processes that highly rely on statistics or for identifying rare reaction channels.
footnotetext: ${}^{a}$ J. R. Macdonald Laboratory, Department of Physics, Kansas State University, Manhattan, KS 66506, USA. E-mail: yijueding@gmail.com
@@ 1 Introduction
In particular, CEI is a powerful technique that is equally sensitive to both light and heavy atoms. CEI retrieves molecular structure by rapidly removing multiple electrons and measuring the asymptotic momenta of ionic fragments. Thanks to decades of development, CEI can be applied to molecules consisting of a couple of atoms to over ten atoms in a molecule10, 11, 12, 13. In a time-resolved manner, CEI has been used to identify isomerization processes14 or isolate a rare roaming channel15.
Methyl iodide ($\text{CH}_{3}\text{I}$) is a prototype molecule to study photochemical reaction dynamics due to its simplicity and rich variety of photochemical phenomena, and thereby has attracted research interests for a long time16, 17, 18, 19, 20, 21. CEI experiments have been conducted on this molecule both near the ground state equilibrium22, 23 and during photodissociation in a time-resolved manner24, 25, 26, 27. One intriguing feature is the nondiabatic dynamics during photodissociation owing to a conical intersection (CI) induced by the strong spin-orbit coupling effect. Non-adiabtic transitions near such a CI result in two dissociation channels corresponding to $\text{I}(^{2}\text{P}_{3/2})$ and $\text{I}^{*}(^{2}\text{P}_{1/2})$ products, respectively.
In the theoretical side, several trajectory calculations have been performed on potential energy surfaces (PES) of $\text{CH}_{3}\text{I}$ excited states over time28, 29, 30, as well as ab initio molecular dynamics (AIMD) calculations31, 32 and quantum mechanical calculations33, 34, 35 that have been conducted to study the A-band photodissociation. However, accurate simulations of both the photodissociation and the Coulomb explosion process to quantitatively simulate the observvables from the pump-probe experiments have not been performed so far. Particularly, no molecular dynamics simulations have been conducted for the doubly charged ionic state, which is important for the Coulomb explosion process.
In this work, we study the UV induced dissociation of $\text{CH}_{3}\text{I}$ probed by CEI in a time-resolved manner through trajectory calculations on ab initio PESs of both the neutral molecule and the doubly charged cation. Since CEI is a statistical method that relies on a large number of events, theoretical methods like AIMD that calculate electronic structures on-the-fly are of limited use as they can typically only predicted a few exemplary trajectories due to high computational cost. To address this issue, we develop an efficient program that performs fast trajectory calculations on both the neutral and ionic surfaces with the inclusion of non-adiabatic effects.
@@ 2.1 Ab initio electronic structure calculation
All electronic structures are computed at the multi-reference configuration interaction (MRCI)36, 37 level of theory. The active space is 6 electrons in 5 orbitals for both the neutral $\text{CH}_{3}\text{I}$ molecule and the $\text{CH}_{3}\text{I}^{2+}$ cation. The active space includes two nearly degenerate lone pair ($n$) orbitals of the I atom, the C-I bonding ($\sigma$) and antibonding ($\sigma^{*}$) orbitals, for the purpose of accurately describing the A-band excited states that are dominated by $n\rightarrow\sigma^{*}$ transitions. One additional orbital is added to the active space to improve convergence.
The state-averaged multi-configuration self-consistent field (SA-MCSCF)38, 39, 40 calculation is first carried out to obtain the reference wavefunctions. For the neutral $\text{CH}_{3}\text{I}$ molecule, 3 lowest singlet states and 3 lowest triplet states are computed and optimized with equal weights in the SA-MCSCF procedure to cover all A-band excited states. For the $\text{CH}_{3}\text{I}^{2+}$ cation, 7 singlets and 4 triplets are computed and optimized with equal weights in the SA-MCSCF procedure to cover all possible atomic terms of the $\text{I}^{+}$ cation. The MRCI calculation is then performed by adding single and double excitations to external configurations from the MCSCF reference wavefunctions.
@@
The Dunning type correlation-consistent basis sets41 are employed in our electronic structure calculations. Specifically, we use cc-pVTZ basis sets for hydrogen and carbon atoms. We also use cc-pVTZ-PP basis set for the iodine atom, where the inner 28 core electrons are described by a relativistic pseudo-potential42, 43.
All ab initio calculations are performed using the MOLPRO quantum chemistry package44, 45.
@@ 2.2 Analytic potential energy surfaces
In order to obtain an accurate description of the electronic energies and the corresponding gradients, we calculate over 40,000 geometries for $\text{CH}_{3}\text{I}$ and over 20,000 geometries for $\text{CH}_{3}\text{I}^{2+}$. The PES is constructed under $\{R,\alpha,\theta\}$ degrees of freedom, where R is the C-I distance, $\alpha$ is the angle between the C-H vector and the $C_{3}$ axis of symmetry, and $\theta$ is the angle between the C-I vector and the $C_{3}$ axis of symmetry. Other degrees of freedom are kept at equilibrium geometry. Particularly, ab initio electronic structure calculations are performed on a very dense grid ($\Delta R=0.01\text{\AA}$) near the conical intersection between ${}^{3}\text{Q}_{0}$ and ${}^{1}\text{Q}_{1}$ states in order to accurately characterize the nonadiabatic effects. Analytic representations of the adiabatic PESs are generated using a multi-dimensional cubic spline interpolation algorithm. Because the grid is sufficiently dense, such an direct interpolation is able to characterize the conical intersection precisely.
The ab initio potential energies are calculated up to $R_{c}=5$ for $\text{CH}_{3}\text{I}$ and up to $R^{ion}_{c}=10$ for $\text{CH}_{3}\text{I}^{2+}$, beyond which the energy gradient of $\text{CH}_{3}\text{I}$ is extremely small($|dV_{\text{CH}_{3}\text{I}}/dR|<0.01$ eV/Å), and the potential of $\text{CH}_{3}\text{I}^{2+}$ is almost purely Coulombic($|d(V_{\text{CH}_{3}\text{I}^{2+}}-1/R)/dR|<0.01$ eV/Å). In order to simulate the the pump-probe process at large delays (up to 500 fs) and to obtain the asymptotic momenta of the ionic fragments after Coulomb explosion, it is natural to extend the PES to infinite C-I distance.
The interaction between the neutral I and $\text{CH}_{3}$ fragments is dominated by the van der Waals force at large internuclear distances, which is in the order of $R^{-6}$. This high-order interaction is negligible, and therefore the $\text{CH}_{3}\text{I}$ potential beyond $R_{c}$ is written as
$V_{\text{CH}_{3}\text{I}}\approx V_{CH_{3}}+V_{I}\qquad(R>R_{c}),$ (1)
@@
The interaction between the ionic $\text{I}^{+}$ and $\text{CH}_{3}^{+}$ fragments is dominated by the Coulomb force at large internuclear distances. We also include a $1/R^{4}$ term to describe the charge-dipole interaction. Thus, the $\text{CH}_{3}\text{I}^{2+}$ potential beyond $R^{ion}_{c}$ is written as
$V_{\text{CH}_{3}\text{I}^{2+}}\approx V_{\text{CH}_{3}^{+}}+V_{\text{I}^{+}}+(\frac{1}{R}-\frac{C_{4}}{R^{4}})\qquad(R>R^{ion}_{c}),$ (2)
@@ 2.3 Photodissociation-Coulomb explosion trajectory calculation
In order to directly simulate the observables from time-resolved CEI experiments, we develop a molecular dynamics program to simulate photodissociation followed by Coulomb explosion in a single trajectory calculation on pre-built analytic PESs of $\text{CH}_{3}\text{I}$ and $\text{CH}_{3}\text{I}^{2+}$. At the first step, the initial conditions are sampled from a classical Boltzmann distribution of normal mode coordinates of the ground electronic state at a certain effective temperature. For UV photon absorption near 266 nm, transition to the ${}^{3}\text{Q}_{0}$ state (mostly parallel) is much stronger than to other states in the A-band20, 46. The momenta and positions of each atom do not change during the UV absorption (vertical transition). Thus, similar to other AIMD simulations performed on this molecule32, 31, 26, we start all trajectories on the ${}^{3}\text{Q}_{0}$ electronic state with initial conditions sampled on the ground electronic state. In our simulation, the equation of motion (EOM) is described by the Euler-Lagrange equation $\frac{d}{dt}\frac{\partial L}{\partial\dot{\textbf{q}}}-\frac{\partial L}{\partial\textbf{q}}=0$, where $\textbf{q}=\{\Psi,\Theta,\Phi,R,\alpha,\theta\}$. The Euler angles $\{\Psi,\Theta,\Phi\}$ are included in the EOM to describe the overall rotation of the entire molecule. The molecule is initially prepared at its rotational ground state with random orientation. To propagate the trajectory, the EOM is integrated numerically using an adaptive Runge-Kutta algorithm with a desired relative error of $10^{-7}$. Non-adiabatic effects during the propagation are taken into account using a Landau-Zener surface hopping algorithm, which will be discussed in the following subsection. When the trajectory is calculated up to a pre-specified target time, the probe laser pulse is applied to doubly ionize the molecule and induces the Coulomb explosion. We refer to this target time as the pump-probe delay. The pulse duration and its effects on the nuclear dynamics are not considered in our simulation. So the molecular dynamics are always conducted on field-free PESs. The photoionization event itself is assumed to occur instantaneously on the timescales considered in our simulation. Therefore, the trajectory instantaneously hops to the $\text{CH}_{3}\text{I}^{2+}$ PES and continues to propagate until the Coulomb energy between ionic fragments is below 0.01 eV. Surface hoppings are not considered on the $\text{CH}_{3}\text{I}^{2+}$ PES. We call this type of trajectories photodissociation-Coulomb explosion (PD-CE) trajectories.
@@
Since CEI is highly dependent on statistics, and one trajectory corresponds to one single event in the coincident ion momentum measurement, a large number of trajectories are required to obtain a statistically reliable interpretation of the CEI results. In our simulation, more than 40,000 PD-CE trajectories are calculated in total, specifically, about 20,000 in the asymptotic regime and about 20,000 for short pump-probe delays.
@@ 2.4 Landau-Zener surface hopping algorithm
In the present study, we employ a Landau-Zener model to treat the non-adiabatic effects in the trajectory calculation. The Landau-Zener surface hopping (LZSH) algorithm has been elaborated in previous studies, and was shown to yields reliable results47, 48, 49, 50. Although Tully’s fewest-switch surface hopping (FSSH) algorithm51 is commonly used for modeling non-adiabatic dynamics in a classical manner, it cannot be rigorously derived from quantum mechanics, and thus the time-evolution of the electron wave function is inconsistent with the classical motion of nuclei. One advantage of LZSH is that it can be readily implemented on adiabatic PESs. Moreover, frustrated hops are less likely to occur in the Landau-Zener model.
The Landau-Zener non-adiabatic transition probability in the adiabatic picture is written as52
@@
where $\textbf{q}(t)$ is the nuclear coordinate vector as a function of time, and $Z_{ij}=|V_{i}-V_{j}|$ is the absolute value of the potential energy difference between $i$ and $j$ adiabatic states. During the trajectory propagation, the value of $Z_{ij}$ is monitored at each time step. When a local minimum of $Z_{ij}$ as a function of time is attained, the non-adiabatic transition probability is calculated using Eq. (3). Then, a random number is generated from a uniform distribution between 0 and 1, and is compared with the probability $P_{i\rightarrow j}$ to determine whether the trajectory should hop to a different adiabatic state. If a hop is confirmed, the velocity of the system is rescaled by a factor $\eta$ to maintain the conservation of energy. This requirement can be written as
$\frac{1}{2}\dot{\textbf{q}}^{T}\textbf{M}\dot{\textbf{q}}+V_{i}(\textbf{q})=\frac{1}{2}\eta\dot{\textbf{q}}^{T}\textbf{M}\eta\dot{\textbf{q}}+V_{j}(\textbf{q}),$ (4)

### v2

@@
Yijue Dinga, Loren Greenmana and Daniel Rollesa
[Uncaptioned image] We present a highly efficient approach to directly and reliably simulate the photodissociation followed by Coulomb explosion of methyl iodide. In order to achieve statistical reliability, more than 40,000 trajectories are calculated on accurate potential energy surfaces of both the neutral molecule and the doubly charged cation. Non-adiabatic effects during photodissociation are treated using a Landau-Zener surface hopping algorithm. The simulation is performed analogous to a recent pump-probe experiment using coincident ion momentum imaging [Ziaee et al., Phys. Chem. Chem. Phys. 2023, 25, $9$]. At large pump-probe delays, the simulated delay-dependent kinetic energy release signals show overall good agreement with the experiment, with two major dissociation channels leading to $\text{I}(^{2}\text{P}_{3/2})$ and $\text{I}^{*}(^{2}\text{P}_{1/2})$ products. At short pump-probe delays, the simulated kinetic energy release differs significantly from the values obtained by a purely Coulombic approximation or a one-dimensional description of the dicationic potential energy surfaces, and shows a clear bifurcation near 12 fs, owing to non-adiabatic transitions through a conical intersection. The proposed approach is particularly suitable and efficient in simulating processes that highly rely on statistics or for identifying rare reaction channels.
footnotetext: a J. R. Macdonald Laboratory, Department of Physics, Kansas State University, Manhattan, KS 66506, USA. E-mail: yijueding@gmail.com
@@ 1 Introduction
In particular, CEI is a powerful technique that is equally sensitive to both light and heavy atoms. CEI retrieves molecular structure by rapidly removing multiple electrons and measuring the asymptotic momenta of ionic fragments. Thanks to decades of development, CEI can be applied to molecules consisting of a couple of atoms to over ten atoms in a molecule10, 11, 12, 13, 14. In a time-resolved manner, CEI has been used to identify isomerization processes15, isolate a rare roaming channel16, or study the Jahn-Teller effect17, 18.
Methyl iodide ($\text{CH}_{3}\text{I}$) is a prototype molecule to study photochemical reaction dynamics due to its simplicity and rich variety of photochemical phenomena, and thereby has attracted research interests for a long time19, 20, 21, 22, 23, 24. CEI experiments have been conducted on this molecule both near the ground state equilibrium25, 26 and during photodissociation in a time-resolved manner27, 28, 29, 30. One intriguing feature is the nondiabatic dynamics during photodissociation owing to a conical intersection (CI) induced by the strong spin-orbit coupling effect. Non-adiabtic transitions near such a CI result in two dissociation channels corresponding to $\text{I}(^{2}\text{P}_{3/2})$ and $\text{I}^{*}(^{2}\text{P}_{1/2})$ products, respectively.
In the theoretical side, several trajectory calculations have been performed on potential energy surfaces (PES) of $\text{CH}_{3}\text{I}$ excited states over time31, 32, 33, as well as ab initio molecular dynamics (AIMD) calculations34, 35 and quantum mechanical calculations36, 37, 38 that have been conducted to study the A-band photodissociation. However, accurate simulations of both the photodissociation and the Coulomb explosion process to quantitatively simulate the observables from the pump-probe experiments have not been performed so far. Particularly, no molecular dynamics simulations have been conducted for the doubly charged ionic state, which is important for the Coulomb explosion process.
In this work, we study the UV-induced dissociation of $\text{CH}_{3}\text{I}$ probed by CEI in a time-resolved manner through trajectory calculations on ab initio PESs of both the neutral molecule and the doubly charged cation. Since CEI is a statistical method that relies on a large number of events, theoretical methods like AIMD, which calculate electronic structures on-the-fly, may not be cost effective to run for a sufficient number of trajectories to obtain proper statistics. To address this issue, we develop an efficient program that performs fast trajectory calculations on both the neutral and ionic surfaces with the inclusion of non-adiabatic effects.
@@ 2.1 Ab initio electronic structure calculation
All electronic structures are computed at the multi-reference configuration interaction (MRCI)39, 40 level of theory. The active space is 6 electrons in 5 orbitals for both the neutral $\text{CH}_{3}\text{I}$ molecule and the $\text{CH}_{3}\text{I}^{2+}$ cation. The active space includes two nearly degenerate lone pair ($n$) orbitals of the I atom, the C-I bonding ($\sigma$) and antibonding ($\sigma^{*}$) orbitals, for the purpose of accurately describing the A-band excited states that are dominated by $n\rightarrow\sigma^{*}$ transitions. Besides these four orbitals, one nearly empty orbital above the $\sigma^{*}$ orbital is added to the active space of $\text{CH}_{3}\text{I}$ and one nearly doubly-occupied orbital below the $\sigma$ orbital is added to the active space of $\text{CH}_{3}\text{I}^{2+}$ to achieve better convergence of energy.
The state-averaged multi-configuration self-consistent field (SA-MCSCF)41, 42, 43 calculation is first carried out to obtain the reference wavefunctions. For the neutral $\text{CH}_{3}\text{I}$ molecule, 3 lowest singlet states and 3 lowest triplet states are computed and optimized with equal weights in the SA-MCSCF procedure to cover all A-band excited states. For the $\text{CH}_{3}\text{I}^{2+}$ cation, 7 singlets and 4 triplets are computed and optimized with equal weights in the SA-MCSCF procedure to cover all possible atomic terms of the $\text{I}^{+}$ cation. The MRCI calculation is then performed by adding single and double excitations to external configurations from the MCSCF reference wavefunctions.
@@
The Dunning type correlation-consistent basis sets44 are employed in our electronic structure calculations. Specifically, we use cc-pVTZ basis sets for hydrogen and carbon atoms. We also use cc-pVTZ-PP basis set for the iodine atom, where the inner 28 core electrons are described by a relativistic pseudo-potential45, 46.
All ab initio calculations are performed using the MOLPRO quantum chemistry package47, 48.
@@ 2.2 Analytic potential energy surfaces
In order to obtain an accurate description of the electronic energies and the corresponding gradients, we calculate about 40,000 geometries for $\text{CH}_{3}\text{I}$ and about 20,000 geometries for $\text{CH}_{3}\text{I}^{2+}$. The PES is constructed under $\{R,\alpha,\theta\}$ degrees of freedom, where R is the C-I distance, $\alpha$ is the angle between the C-H vector and the $C_{3}$ axis of the $\text{CH}_{3}$ component, and $\theta$ is the angle between the C-I vector and the $C_{3}$ axis of the $\text{CH}_{3}$ component. The umbrella motion (characterized by $\alpha$) and the C-I bending (characterized by $\theta$) are strongly coupled with the C-I bond stretching during dissociation. The C-I bending motion also plays an important role in non-adiabatic transitions. Particularly, ab initio electronic structure calculations are performed on a very dense grid ($\Delta R=0.01\text{\AA}$) near the conical intersection between ${}^{3}\text{Q}_{0}$ and ${}^{1}\text{Q}_{1}$ states in order to accurately characterize the nonadiabatic effects. Analytic representations of the adiabatic PESs are generated using a multivariate cubic spline interpolation algorithm. Because the grid is sufficiently dense, such a direct interpolation is able to characterize the conical intersection with high fidelity.
The ab initio potential energies are calculated up to $R_{c}=5$ for $\text{CH}_{3}\text{I}$ and up to $R^{\mathrm{ion}}_{c}=10$ for $\text{CH}_{3}\text{I}^{2+}$, beyond which the energy gradient of $\text{CH}_{3}\text{I}$ is small($|dV_{\text{CH}_{3}\text{I}}/dR|<0.01$ eV/Å), and the potential of $\text{CH}_{3}\text{I}^{2+}$ is almost purely Coulombic($|d(V_{\text{CH}_{3}\text{I}^{2+}}-1/R)/dR|<0.01$ eV/Å). In order to simulate the the pump-probe process at large delays (up to 500 fs) and to obtain the asymptotic momenta of the ionic fragments after Coulomb explosion, it is natural to extend the PES to infinite C-I distance.
The interaction between the neutral I and $\text{CH}_{3}$ fragments is dominated by the van der Waals force at large internuclear distances, which is in the order of $R^{-6}$. We neglect this high-order interaction, and therefore the $\text{CH}_{3}\text{I}$ potential beyond $R_{c}$ is written as
$V_{\text{CH}_{3}\text{I}}\approx V_{\text{CH}_{3}}+V_{\text{I}}\qquad(R>R_{c}),$ (1)
@@
The interaction between the ionic $\text{I}^{+}$ and $\text{CH}_{3}^{+}$ fragments is dominated by the Coulomb force at large internuclear distances. We also include a $1/R^{4}$ term to describe the charge-dipole interaction49. This isotropic term characterizes both the charge-permanent dipole interaction after averaging the dipole orientation and the charge-induced dipole interaction. Thus, the $\text{CH}_{3}\text{I}^{2+}$ potential beyond $R^{\mathrm{ion}}_{c}$ is written as
$V_{\text{CH}_{3}\text{I}^{2+}}\approx V_{\text{CH}_{3}^{+}}+V_{\text{I}^{+}}+(\frac{1}{R}-\frac{C_{4}}{R^{4}})\qquad(R>R^{\mathrm{ion}}_{c}),$ (2)
@@ 2.3 Photodissociation-Coulomb explosion trajectory calculation
In order to directly simulate the observables from time-resolved CEI experiments, we develop a molecular dynamics program to simulate photodissociation followed by Coulomb explosion in a single trajectory calculation on pre-built analytic PESs of $\text{CH}_{3}\text{I}$ and $\text{CH}_{3}\text{I}^{2+}$.
At the first step, the initial conditions are sampled from a classical Boltzmann distribution [^2] Zero-point energy (ZPE) is not included in this sampling method, so there is no ZPE leakage issue in the trajectory calculation. The ZPE leakage issue would not be significant even if ZPE were considered since the high-frequency CH stretching modes are absent due to reduced dimensionality. The agreement between our simulation and the experiment also justifies our assumption. of normal mode coordinates of the ground electronic state at effective temperature $T=100\sim 200$ K. Since the PESs are constructed in reduced dimension, the normal mode coordinates $\eta_{i}(i=1,2,3)$ can be written as $\{\eta_{1},\eta_{2},\eta_{3}\}=\textbf{U}\{R,\alpha,\theta\}$, where U is the transformation matrix. The Boltzmann probability distribution is given by
$P(\eta_{i},\dot{\eta}_{i})\propto\exp\{-(m_{i}\dot{\eta}_{i}^{2}+m_{i}\omega_{i}^{2}\eta_{i}^{2})/2k_{B}T\}\quad(i=1,2,3)$ (3)
where $k_{B}$ is the Boltzmann constant, $m_{i}$ is the generalized mass for each normal mode coordinate, and $\omega_{i}$ is the corresponding normal mode frequency. The normal mode analysis on the ground state PES reveals $\{\omega_{1},\omega_{2},\omega_{3}\}=\{530,880,1308\}\text{ cm}^{-1}$, corresponding to C-I stretching, C-I bending, and $\text{CH}_{3}$ umbrella motion, respectively. These frequencies are in overall agreement with the reported experimental values {533,882,1251} $\{533,882,1251\}\text{ cm}^{-1}$50, 51, 52. The remaining internal degrees of freedom are kept the same as in the equilibrium geometry.
@@
For UV photon absorption near 266 nm, transition to the ${}^{3}\text{Q}_{0}$ state (mostly parallel) is much stronger than to other states in the A-band23, 53. The momenta and positions of each atom do not change during the UV absorption (vertical transition). Thus, similar to other AIMD simulations performed on this molecule35, 34, 29, we start all trajectories on the ${}^{3}\text{Q}_{0}$ electronic state with initial conditions sampled on the ground electronic state. In our simulation, the equation of motion (EOM) is described by the Euler-Lagrange equation $\frac{d}{dt}\frac{\partial L}{\partial\dot{\textbf{q}}}-\frac{\partial L}{\partial\textbf{q}}=0$, where $\textbf{q}=\{\Psi,\Theta,\Phi,R,\alpha,\theta\}$. The Euler angles $\{\Psi,\Theta,\Phi\}$ are included in the EOM to describe the overall rotation of the entire molecule, which is likely to affect the rotational states of the resulting fragments after photodissociation and Coulomb explosion. The molecule is initially prepared at its rotational ground state with random orientation. To propagate the trajectory, the EOM is integrated numerically using an adaptive Runge-Kutta algorithm with a desired relative error of $10^{-7}$. Non-adiabatic effects during the propagation are taken into account using a Landau-Zener surface hopping algorithm, which will be discussed in the following subsection. When the trajectory is calculated up to a pre-specified target time, the probe laser pulse is applied to doubly ionize the molecule and induces the Coulomb explosion. We refer to this target time as the pump-probe delay. The pulse duration and its effects on the nuclear dynamics are not considered in our simulation. So the molecular dynamics are always conducted on field-free PESs. The photoionization event itself is assumed to occur instantaneously on the timescales considered in our simulation. Therefore, the trajectory instantaneously hops to the $\text{CH}_{3}\text{I}^{2+}$ PES and continues to propagate until the Coulomb energy between ionic fragments is below 0.01 eV. To simplify the problem, non-adiabatic effects are not considered at the Coulomb explosion stage. We call this type of trajectories photodissociation-Coulomb explosion (PD-CE) trajectories.
@@
Since CEI is highly dependent on statistics, a large number of trajectories are required to obtain a statistically reliable interpretation of the CEI results. In our simulation, more than 40,000 PD-CE trajectories are calculated in total, specifically, about 20,000 in the asymptotic regime and about 20,000 for short pump-probe delays.
@@ 2.4 Landau-Zener surface hopping algorithm
In the present study, we employ a Landau-Zener model to treat the non-adiabatic effects in the trajectory calculation. The Landau-Zener surface hopping (LZSH) algorithm has been elaborated in previous studies, and was shown to yields reliable results54, 55, 56, 57. The LZSH algorithm is easier to implement compared to the commonly used fewest-switch surface hopping (FSSH) algorithm58, while still yielding similar non-adiabatic results such as the time-dependent electronic population ratio54, 57.
The Landau-Zener non-adiabatic transition probability in the adiabatic picture is written as59
@@
where $\textbf{q}(t)$ is the nuclear coordinate vector as a function of time, and $Z_{ij}=|V_{i}-V_{j}|$ is the absolute value of the potential energy difference between $i$ and $j$ adiabatic states. During the trajectory propagation, the value of $Z_{ij}$ is monitored at each time step. When a local minimum of $Z_{ij}$ as a function of time is attained, the non-adiabatic transition probability is calculated using Eq. (4). Then, a random number is generated from a uniform distribution between 0 and 1, and is compared with the probability $P_{i\rightarrow j}$ to determine whether the trajectory should hop to a different adiabatic state. If a hop is confirmed, the velocity of the system is rescaled by a factor $\gamma$ to maintain the conservation of energy. This requirement can be written as
$\frac{1}{2}\dot{\textbf{q}}^{T}\textbf{M}\dot{\textbf{q}}+V_{i}(\textbf{q})=\frac{1}{2}\gamma\dot{\textbf{q}}^{T}\textbf{M}\gamma\dot{\textbf{q}}+V_{j}(\textbf{q}),$ (5)
@@
It should be noted that in this model a surface hopping decision can be made even when $Z_{ij}$ is large, but in our simulation all trajectories that undergo non-adiabatic transitions show $Z_{ij}<0.2$ eV when the surface hoppings take place. Furthermore, non-adiabatic effects are only considered within 2.2 Å<$2.2\text{ \AA}<R<2.6\text{ \AA}$.